# Baillon-Bruck-Reich revisited: divergent-series parameters and strong convergence in the linear case

### v1

@@ Abstract
The Krasnoselskii-Mann iteration is an important algorithm in optimization and variational analysis for finding fixed points of nonexpansive mappings. In the general case, it produces a sequence converging weakly to a fixed point provided the parameter sequence satisfies a divergent-series condition.
@@
Keywords: Baillon-Bruck-Reich theorem, fixed point, Hilbert space, Krasnoselskii-Mann iteration, nonexpansive mapping.
@@ 1 Introduction
Many problems in optimization and variational analysis can be reduced to finding a point in $F$ (see, e.g., [2], [5], [9], and the references therein). A popular algorithmic approach is to employ the Krasnoselskii-Mann iteration for which we now state a basic convergence result:
@@
We refer the reader also to the recent book [6], dedicated entirely to Krasnoselskii-Mann iterations. When $T$ is linear, the Krasnoselskii-Mann iteration produces a strongly convergent sequence provided that the parameter sequence $(\lambda_{n})_{n\in{\mathbb{N}}}$ is constant:
@@
The remainder of the paper is organized as follows. In Section 2, we present the main result (Theorem 2.2) and its proof. The paper is concluded with various remarks.
@@ 2 Main result
$T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{\varepsilon}=T_{\lambda_{n}}\cdots T_{\lambda_{0}}(\operatorname{Id}-T)v_{\varepsilon}=T_{\lambda_{n}}\cdots T_{\lambda_{0}}v_{\varepsilon}-TT_{\lambda_{n}}\cdots T_{\lambda_{0}}v_{\varepsilon}\to 0.$ (16)
Thus
@@
We conclude with some remarks concering Theorem $2.$
@@
Thanks to an example constructed by Genel and Lindenstrauss (see [7]), it is known that the sequence $(x_{n})_{n\in{\mathbb{N}}}$ generated by the Krasnoselskii-Mann iteration may fail to converge strongly even when $\lambda_{n}\equiv\tfrac{1}{2}$. A posteriori, the underlying nonexpansive mapping cannot be linear.
@@
Suppose that $x_{0}\in X\smallsetminus\{0\}$, and consider the linear nonexpansive isometry $T=-\operatorname{Id}$. Then $F=\{0\}$. Assume that the parameter sequence $(\lambda_{n})_{n\in{\mathbb{N}}}$ lies in $\left[0,\tfrac{1}{2}\right[$. Then the sequence $(x_{n})_{n\in{\mathbb{N}}}$ generated by the Krasnoselskii-Mann iteration satisfies
@@
moreover, $x_{n}\to 0$ $\Leftrightarrow$ $\sum_{n\in{\mathbb{N}}}\lambda_{k}=\infty$. Therefore, if $\sum_{n\in{\mathbb{N}}}(1-\lambda_{n})\lambda_{n}<\infty$, then $(x_{n})_{n\in{\mathbb{N}}}$ does not converge to a point in $F$.
@@
Remark 2.7 (an open problem).
The original paper by Baillon, Bruck, and Reich [1] was situated in a more general setting with respect to the underlying space and the nonexpansive map; in fact, $T$ was merely assumed to be odd rather than linear. At present, we do not know if it is possible to extend Theorem 2.2 from linear to odd nonexpansive mappings.
Acknowledgment
The research of SB was partially supported by a collaboration grant for mathematicians of the Simons Foundation. The research of HHB was partially supported by a Discovery Grant of the Natural Sciences and Engineering Research Council of Canada.

### v2

@@ Abstract
The Krasnosel’skiĭ-Mann iteration is an important algorithm in optimization and variational analysis for finding fixed points of nonexpansive mappings. In the general case, it produces a sequence converging weakly to a fixed point provided the parameter sequence satisfies a divergent-series condition.
@@
Keywords: Baillon-Bruck-Reich theorem, fixed point, Hilbert space, Krasnosel’skiĭ-Mann iteration, nonexpansive mapping.
@@ 1 Introduction
Many problems in optimization and variational analysis can be reduced to finding a point in $F$ (see, e.g., [2], [6], [10], and the references therein). A popular algorithmic approach is to employ the Krasnosel’skiĭ-Mann iteration for which we now state a basic convergence result:
@@
We refer the reader also to the recent book [7], dedicated entirely to Krasnosel’skiĭ-Mann iterations. When $T$ is linear, the Krasnosel’skiĭ-Mann iteration produces a strongly convergent sequence provided that the parameter sequence $(\lambda_{n})_{n\in{\mathbb{N}}}$ is constant:
@@
The remainder of the paper is organized as follows. In Section 2, we present the main result (Theorem 2.3), its proof, and corresponding comments. In Section 3, we offer some concluding remarks.
@@ 2 Main result
Remark $2.$
We are grateful to a reviewer who outlined the following alternative proof of Section 2: Let $A\colon X\rightrightarrows X$ be a maximally monotone linear relation. Combining [4, Fact 2.1.(v) and Theorem 3.2] yields $(\ker A)^{\perp}=\overline{\operatorname{ran}}\,A^{*}=\overline{\operatorname{ran}}\,A$. Then Section 2 follows when one sets $A:=\operatorname{Id}-T$.
@@
$T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{\varepsilon}=T_{\lambda_{n}}\cdots T_{\lambda_{0}}(\operatorname{Id}-T)v_{\varepsilon}=T_{\lambda_{n}}\cdots T_{\lambda_{0}}v_{\varepsilon}-TT_{\lambda_{n}}\cdots T_{\lambda_{0}}v_{\varepsilon}\to 0,$ (16)
where we also used the fact that $T$ commutes with each $T_{\lambda_{i}}$. Thus
@@
We conclude with some remarks concerning Theorem 2.3.
@@
Thanks to an example constructed by Genel and Lindenstrauss (see [8]), it is known that the sequence $(x_{n})_{n\in{\mathbb{N}}}$ generated by the Krasnosel’skiĭ-Mann iteration may fail to converge strongly even when $\lambda_{n}\equiv\tfrac{1}{2}$. A posteriori, the underlying nonexpansive mapping cannot be linear.
@@
Suppose that $x_{0}\in X\smallsetminus\{0\}$, and consider the linear nonexpansive isometry $T=-\operatorname{Id}$. Then $F=\{0\}$. Assume that the parameter sequence $(\lambda_{n})_{n\in{\mathbb{N}}}$ lies in $\left[0,\tfrac{1}{2}\right[$. Using mathematical induction, one checks that the sequence $(x_{n})_{n\in{\mathbb{N}}}$ generated by the Krasnosel’skiĭ-Mann iteration satisfies
@@
moreover, $x_{n}\to 0$ $\Leftrightarrow$ $\sum_{n\in{\mathbb{N}}}\lambda_{n}=\infty$. Therefore, if $\sum_{n\in{\mathbb{N}}}(1-\lambda_{n})\lambda_{n}<\infty$, then $(x_{n})_{n\in{\mathbb{N}}}$ does not converge to a point in $F$.
@@
3 Conclusions
A well-known result by Reich on Krasnosel’skiĭ-Mann iterations yields weak convergence to a fixed point of a nonexpansive mapping $T$ under a general (divergent-series) assumption on the parameters. On the other hand, when $T$ is linear, work by Baillon, Bruck and Reich implies that the convergence is strong but the relaxation parameters needed to be constant. In this paper, we extend the Baillon-Bruck-Reich result from constant to general parameters in the linear case. The limit is explicitly identified, and some limiting examples and extensions are discussed. It would be interesting to know if this result remains true if $T$ is only assumed to be odd instead of linear as was the case in [1].
Acknowledgments
The authors thank the reviewers and the editor for constructive comments which significantly improved the manuscript. The research of SB was partially supported by a collaboration grant for mathematicians of the Simons Foundation. The research of HHB was partially supported by a Discovery Grant of the Natural Sciences and Engineering Research Council of Canada.